OpenStudy (jakelee):

6+x/2=4 two step equation

OpenStudy (texaschic101):

6 + x/2 = 4 --- subtract 6 from both sides x/2 = 4 - 6 x/2 = -2 -- multiply both sides by 2 x = -2 * 2 x = -4 check.. 6 + x/2 = 4 6 + (-4/2) = 4 6 - 2 = 4 4 = 4 (correct) so x = -4